OpenStudy (anonymous):

Find the x-coordinate of the point(s) of inflection on the graph of y = x^3 – 6x^2 + 9x + 3. a. x = 0 b. x = 1 c. x = 2 d. x = 3

OpenStudy (anonymous):

the critical pts are 1 and 3

OpenStudy (anonymous):

@ganeshie8 @Directrix @mathmate @AuroraB @iamawesome1 @Abhisar @wio @iGreen @nincompoop @Preetha @uri @TheSmartOne

OpenStudy (anonymous):

@Michele_Laino @AlexandervonHumboldt2 @Abhisar

OpenStudy (anonymous):

Do I find the second derivative?

OpenStudy (anonymous):

f'(x)= 3x^2-12x+9 f''(x)= 6x-12

OpenStudy (anonymous):

If I set 6x=12 then x=2

OpenStudy (anonymous):

\[find \frac{ d^2y }{ dx^2 }~at~x=1,x=3\] where \[\frac{ d^2y }{ dx^2 }=0\] that point is point of inflexion

OpenStudy (anonymous):

from wht u told me to do i got (1,-6)(3,6)

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

so now what do i do?

OpenStudy (anonymous):

at x=2, f"(x)=0 so point of inflexion is x=2 1. at point of inflexion concavity or convexity changes

OpenStudy (anonymous):



OpenStudy (anonymous):

so the answers 2 then?

OpenStudy (anonymous):

@paki @ganeshie8 @Directrix @mathmate @AuroraB @iamawesome1 @wio @iGreen @nincompoop @Preetha @uri @TheSmartOne @surjithayer @Abhisar @Alexistheamazing @AlexandervonHumboldt2 @Quan99 @wio @Loser66 @Compassionate @uri